Differences between revisions 491 and 492
Revision 491 as of 2019-08-16 13:19:01
Size: 30485
Comment:
Revision 492 as of 2019-08-16 13:43:37
Size: 30761
Comment:
Deletions are marked like this. Additions are marked like this.
Line 11: Line 11:
 * The BS optic has a large peak at 150 mHz.
 * The other suspensions
 * The BS optic has a large peak at 150 mHz (klog [[http://klog.icrr.u-tokyo.ac.jp/osl/?r=9991|9991]]).

==== Plan for the week 19-08-2019 ====
 * Enomoto-kun requested to extend the oplev control loop down to where the micro-seismic peak is. (Fabian.)
 * Assemsment of the motion of the SR optics around the micro-seimic region. (Due today.)

Scraps of text for incorporation in the VIS Meeting Minutes

KAGRA/Subgroups/VIS/MeetingNAOJ

Type B

Report for the week 15-08-2019 (Written by Fabian)

  • Shoda-san improved the SR3 optic yaw motion (klog 9949).

    • Improvement in the frequency band from 1 Hz to 2.3 Hz was achieved by modifying the IM-Y filter. The problem was not enough phase margin.
    • There's still a peak at 1.3 Hz.
    • She also modified the IM SENSALIGN matrix (diagonalization according to resonance peaks?)
    • Transfer functions should be measured again.
  • The BS optic has a large peak at 150 mHz (klog 9991).

Plan for the week 19-08-2019

  • Enomoto-kun requested to extend the oplev control loop down to where the micro-seismic peak is. (Fabian.)
  • Assemsment of the motion of the SR optics around the micro-seimic region. (Due today.)

Type B

Report for the week 05-08-2019 (Written by Fabian)

  • After Shoda-san's report (entry 9797) I checked the condition of the filters in FLOAT state.

    • After a few tests I realized the whitening chassis didn't have a whitening filter board. See entry 9844.

    • I replaced the chassis with an appropriate one. See entry 9876.

  • I checked the configuration of the hardware whitening filters by checking the 32 bit dip switch boards for each supension. See entry 9823 and 9876.

    • The switch configuration is the same for SRM and SR2.
    • For BS and SR3 the configurations are different.
    • Additionally, the configuration of the two boards used for the BS oplev are different from each other.
    • This situation should be reviewed.
    • Type Bp and Type A don't seem to be using software anti-whitening filters for oplevs. I haven't checked their hardwate whitening filters' configuration though.
  • At the begining of the week I continued noise hunting in SR3. See entry 9807. The goal was to identify the limiting elements for the control system in ALIGNED state. See entries

    • The motion of the IP-L induces motion in TM-P and TM-L in the frequency band from 0.135 Hz to 0.4 Hz. This should be fixed at the IP. There's no inertial damping yet.

    • The motion of the IM-L induces motion in TM-P nd TM-L in the frequency band from 0.3 Hz to 0.8 Hz. Likely, this should be fixed with IML2TMP and IML2TML control filters.

    • The motion of the IM-P induces motion in TM-P and TM-L in the frequency band from 0.135 Hz to 1 Hz. The motion in the band from 0.135 to 0.4 Hz is likely ground motion coming from the IP, however, the motion from 0.4 Hz to 1 Hz should be fixed with IMP2TML and IMP2TMP control filters.

    • The motion in IM-Y induces motion TM-Y between 1 Hz and 2.3 Hz. This should be fixed with IMY2TMY control filters.

    • I began measuring transfer functions (IM-L to TM-L) but I should finish the setting the software anti-whitening filter first.

Plan for the week 15-08-2019

  • Finish measuring transfer functions.
  • Begin the design of filters.

Report for the week 7/29 (Written by Fabian)

  • I offloaded fishing rods in GAS filters and IPs in BS and SRM.
  • I implemented scripts for the automatic measurement of amplitude spectral densities (ASDs) of suspension sensors (usually used for residual motion estimation).
  • Highlights of recent ASD measurements can be found in entry 9726. The residual motion was measured in ALIGNED and FLOAT states.

    • In SRM and SR3 the motion in ALINED state was smaller than in FLOAT state, which is expected. See entries and respectively .
    • In SR2, the motion in pitch and yaw displacement is larger in ALIGNED state, and the longitudinal velocity is the same in both states.

    • In BS, the longitudinal displacement is the same in both states.

  • Some features of the measured ASDs are the following:
    • In SRM, SR3 and SR2 in pitch there is either a large bump (SRM and SR3) or excitation (SR2) between 0.3 Hz and 1 Hz. 9756.

    • In SRM pitch the control system seems to be introducing noise in two frequency bands, before and after the bump. See 9756.

    • In SR3 pitch the control loops seems to be introducing noise between 1 Hz and 2 Hz. See 9755.

    • In SR2 pitch the control loops seems to be introducing noise between 0.380 Hz and 2 Hz. See 9752.

    • In SR2 yaw the control loops seems to be introducing noise between 0.045 Hz and 0.230 Hz. See 9752.

    • In SR2 longitudinal speed the control system seems to excite a large peak at 0.705 Hz. See 9752.

Plan for the week 8/5

  • Check the BS amplitude spectral densities.
  • Check the longitudinal degree of freedom in SRs.
  • Aim to identify noise sources.

Type B

Report for the week 7/22 (Written by Fabian)

  • Shoda-san set the auto-measurement scripts and directories. She measured the transfer functions having the suspensions in FLOAT state.
  • SR3 IM H2 OSEM channel noise hunting (klog 9613 and 9629).The symptom was that the noise in SR3 IM H2 channel was 4.7 times larger than in similar channels. When the suspension was quiet H2 was not measuring any displacement but only noise. We did several measurements:

    • We removed the cable from the satellite box at the AA filter input and terminated the input: the noise was still there.
    • We terminated the ADC from the inside of the AA filter box: the noise was gone, suggesting the problem was the AA filter.
    • We replaced the AA filter and repeated the measurement with the input terminated: the noise was still there.
    • We disconnected all the cables from the input of the AA filter and terminated the offending channel: the noise was still there.
    • We also changed cables and tried connecting the AA filter to the SRM rack but the noise was still there.
    • The solution was to use other channels for all IM-H OSEMs. The real-time model was changed accordingly.
  • The SR3 LVDT driver broke down (klog 9659).

    • We changed the AA filter power cable and when we turned the filter back on the whole rack turned off momentarily and came back on, with the exception of the LVDT driver.
    • We tried turning it on but it kept switching off by itself. Tomura-san suspects there's a grounding problem.
    • We decided to change the driver.
  • LDVT driver replacement and adjustment (klog 9658 and 9025).

    • For each channel there are three resisitors which require adjustment and determine an input gain, an output gain and a phase.
    • The phase was originally measured with an oscilloscope and the resistor was adjusted using also the oscilloscope.
    • The output gain resistor was set to the desired value using the multi-meter.
    • For the input gain we tried mesuring the signal amplitude with the oscilloscope (per Nikhef instructions) but we didn't see any obvios change when changing the resistence value.
    • Alternatively we measured the input resistence values in the old cards and set the new ones accordingly.
    • We did quick measurements of transfer functions and they looked alright. Measurements with better frequency resolution are needed.

Plan for the week 7/29

  • Noise assessment in SR3.
    • IM-Y integrated RMS displacement with control system on. Compare with klog 9494.

    • Terrence verbally reported LVDTs are particulrly noisy.
    • I would like to measure ASD for all sensors (in FLOAT state) and compare with other suspensions.
    • Miyakawa-san is interested to find the origin of the noise in one particular channel.
  • Check SRM BF position. The set point should be zero (klog 8672) but today I realized it was set to 100 um sometime.

  • Check if we need to offload coil-magnet actuation onto fishing rods.
  • Remove the software de-white (OSEMs) and antiWTN (oplevs) filters used in some or all suspensions. We don't have hardware whitening filters at the input.
  • Compile information about Guardian states and make it available to all.

Vacation/Travel Info

  • Terrence left Kamioka. He will be back on September.

Type B (Mark)

Report for the week 7/15 (Written by Fabian)

  • SR3 IM H2 noise hunting.
  • SRM oplev condition: when the suspension is in ALIGNED state the oplev (pitch) is saturated and yaw within an inconvenient range (non-linear). See klog entry 9562.

    • Currently we don't have a good reference for the optic alignment and we cannot decide by how much we should move the QPD. This situation is not new and we tried to cope with it by using historical data of IM DC feedback signal in a good orientation for the DRMI. However, this strategy didn't work.
    • The consequence is that in ALIGNED state the pitch feedback DC signal drifts.
    • Kokeyama-san and Valera aligned the mirror, provided a reference and we were able to move the QPD. See entry 9579.

    • The big flange was moved to another place.
  • BS IM2TM transfer functions were measured in ALIGNED state with the control system on. See entry 9571.

  • HPCD modification. See entry 9541 and comments.

Plan for the week 7/22

  • Go back to the tunnel on Monday to continue the investigation in SR3 IM H2 OSEM noise.

Vacation/Travel Info

  • Terrence goes back to Hong Kong on 7/27 (sat), last day in Kamioka: 7/25 (Thurs).

Type B (Mark)

Report for the week 7/8 (Written by Fabian)

  • There was DAQ channel update on Wednesday.
  • After making a wind shield for SR3 oplev last week Terrence revisited the design of the IM to TM control filter with oplev and tried to reduce the residual motion peak at 1.4 Hz reported by Izumi-san.
  • Last week we moved all the SRs oplev QPDs because they were off by an unsuitable amount. After the interfrometer was aligned again we realized some QPDs will likely require to be adjusted again (optimal position would be within ±50 in normalized displacement units): SR2 tilt and SRM tilt require further adjustment.
  • During DRMI commissioning people realized the SR3 yaw motion was too large and was compromising the stability.
    • The yaw peak-to-peak motion was 6 µrad and involved IM/RM/TM resonat modes at ~0.147 Hz and ~0.172 Hz.
    • Terrence realized the amplitude decreased just by turning off the IM yaw control. He tried to damp IM-Y with the oplev but more time is required.
    • We realized the H2 IM OSEM is likely too noisy and excites resonant motion. We went to the mine, checked the LED current monitor but couldn't find anything obviously wrong using the oscilloscope. More investigation is needed.
    • Enomoto-kun pointed out that the IM2TM oplev DC control might be marginally stable. Terrence measure the IM-Y to TM-Y transfer function and designed the filter again allowing a suitable phase margin at 0.14 Hz.
  • In the satellite boxes the label "CURR Mon. for PD" should be "CURR Mon. for LED".

  • SR3 and SRM oplev sled are decreasing their power (e.g. 50% in 150 days in SR3). The situation has been reported to Akutsu-san already.
  • Fabian came up with a preliminary plan to search for common mode motion in Type B systems (after talking with Miyo-kun, Terrence and Lucia.)
  • Report the status of the control system.

Plan for the week 15/8

  • Go back to the tunnel on Tuesday to continue the investigation in SR3 IM H2 OSEM noise.
  • Measure and provide IM-L to TM-P transfer functions to the commissioning team (Stefan's request this morning).

Vacation/Travel Info

  • Terrence goes back to Hong Kong on 7/27 (sat), last day in Kamioka: 7/25 (Thurs).

Type B (Mark)

Report for the week 7/1 (Written by Fabian)

  • Fabian and Terrence went through the process of oplev diagonalization for the purposes of knowledge transfer and procedure documentation.
    • There were small corrections to the procedure.
    • Terrence wrote the oplev section of the Type B paper.
    • Fabian began writing a rough draft of the procedure.
  • With the agreement of the comissioning team we decided to move the position of all of oplev QPDs. When the optics were aligned they were away from linear range.
    • Although the badly affected channels were only three, we decided to adjust all of them.
    • The procedure is described in klog entry 9411.

    • In order to keep the aligmment information, we took note of the IM DC control signal necessary to achive the alignment and and off-loaded the actuation onto the TEST channels while the oplev was adjusted with the control system off.
    • In the case of SRM the oplev was to the point of saturation and was not rendering a usable feedabck signal when the optic was alledgely in aligned state. In this case we looked for the appropriate amount of feedback in stored data from a time in which the DRMI had been locked.
  • On Friday we went to the tunnel in order to put a better wind shield around SR3 oplev optics. The aim is to reduce the noise and then improve the performance of the control system.
  • We found large and heavy blank flanges leaning on the large SR3 and SRM frames. Removing them will likely disturb the position of the IPs, which would require readjustig the QPDs again. We would like the flanges to remain there for O3 if possible. Fabian will write a message to the vacuum group with our request.

Plan for the week 7/8

  • Terrence and Fabian will continue writing the oplev diagonalization procedure.
  • Fabian will install the rolled-off the DC control filters.

Type B (Mark)

Report for the week 6/24 (Written by Terrence and Fabian)

  • Type B watchdog/DACKILL is split for tower and payload. Now, tripping the payload watchdog will only trigger the payload DACKILL while tripping the tower watchdog will trigger both watchdogs and DACKILL. The MEDM screen is updated to accommodate the changes. The screen is also modified so the reset buttons for watchdogs and DACKILL are all available on the same screen. In the middle of this work, a bug is discovered in the ModerateWD.c where setting 0 to the ramp down time would cause the WD to loop in an endless ramp down cycle.
  • The Guardian is updated to recognize if the payload is tripped and decide which tripped state it should go to, "TRIPPED" or "PAYLOAD_TRIPPED". So, if both tower and payload controls are on while the payload watchdog tripped, the guardian will jump to the "PAYLOAD_TRIPPED" state and shutdown payload controls only. When the WD and DACKILL are reset, the "TRIPPED" state will dump you out to the safe state where all controls are disengaged, while the "PAYLOAD_TRIPPED" state will dump you out to the "ENGAGING_IM_DAMPING" state where the tower controls are still on.
  • VIS monitor is updated. It will monitor the raw sensor input and actuator output and check if they are within the operation range. The indicator will turn orange when it is 50% out of range and will turn red if it is 90% out of range. The indicator for actuators will turn blue if they are not in use. The limits of the operating range can be set in userapps/vis/k1/guardian/VIS_DICT.py. The range for OL_PIT and OL_YAW are default to be [-1,1], the lower limit of the OL_SUMs are default to be 1000 counts while those for the other are default to be [-32768,32768].
  • If you know the range of certain sensors and actuators, you are strongly encouraged to tweak the thresholds in the file. Otherwise, the function of the monitor will be very limited.

  • Fabian checked again the GAS filter control noise:
    • My previous calculation considered damping filters and those injected too much noise in the observation frequency band. Rolling the filters off with (high order) Butterworth filters rendered the system unstable.
    • I then considered DC control filters only and I was able reduce the injectetd noise to acceptable levels while keeping the loop stable. As pointed out by Shoda-san, this was the strategy devised for the Type B prototype.
    • This has only been tested in software, not in the actual system. I aim to do implement it next week.
    • This strategy will work only if the resonant motion of the system is not excited after disabling the payload and GAS chain damping control. On the 23rd of May Terrence measured the residual motion with the control system (including IP) on and off. There was no subtantial difference between the two. I aim to take more measurements.
  • Terrence and Fabian are currently going through the oplev diagonalization procedure and will write down the procedure.
  • Hirata-san worked in the design of the baffles for the BS.

Plan for the week 7/1

  • Terrence and Fabian will continue writing the oplev diagonalization procedure.
  • Fabian will install the rolled-off the DC control filters.

Vacation/Travel Info

  • Terrence goes back to Hong Kong on 7/27 (sat), last day in Kamioka: 7/25 (Thurs).
  • Mark's final day 6/28.

Type B (Mark)

Report for the week 6/17

  • Izum-san pointed out the SR3 oplev control loop was not able to damp resonances at 1.4 Hz and 1.2 Hz (klog 9290). Terrence checked (klog 9299):

    • He found that the optic TFs were coupled to about 10%. I assume something changed because, as far as I understand (Fabian), the oplev had been succesfully diagonalized before to less than 1% coupling.
    • He moved the IP and compared IP LENGTH with OPLEV LENGTH noting the the oplev was reading only 70% of the IP displacement.
    • In this conditions he diagonalized the oplev but failed. There was still coupling.
    • He introduced a factor to have the OPLEV LENGTH measurement equal to IP LENGTH measurement.
    • He successfully diagoinalized the oplev.
    • He then proceeded to checked the filter but he was unable to reduce the residual motion. Noise is the suspect. Having appropriate wind shield is recommended.
  • Fabian worked in rolling-off the control filters of the GAS filters.
    • According to klog 8979 the GAS filter control noise has to decrease by 5 orders of magnitude in order to meet the requirement at 10 Hz.

    • A Butterworth filter of order 20 with a cut-off frequency of 5 Hz seems to reduce the noise but renders the system unstable.
    • More work is necessary.
  • Fabian joined Enomoto-kun's lecture on how to align the X Fabry Perot cavity with green light.

Plan for the week 6/24

  • Terrence will work on...
  • Fabian will continue working in the rolling-off of the filters of the GAS filters control loops.

Vacation/Travel Info

  • Terrence goes back to Hong Kong on 7/27.
  • Mark's final day 6/28.

Type B (Mark)

Report for the week 6/10

  • Supporting the commisioning team.

BS
  • Type B inertial damping. This means geophone diagonalization, blending filter optimizing, etc.
  • Optimization gives an odd looking blending filter that blends at ~120 Hz, but works fine in IP L. 9199

  • During PRMI lock, BS residual motion was too high which misaligned the beam too much from time to time and making the lock unstable.
  • To fix this, I improved OL windshield so I can study the residual motion of the BS. Then I copied IP L blending filter to IP T and IP Y and cranked up the IP damping and TM damping gain. This suppressed the residual motion so the requirements are met. After this, the BS was not causing any problem to the lock anymore. 92199235

Type B (Mark)

Report for the week 6/3

SRM Inertial Damping and Residual Motion Stuff
  • See Klog 9064

  • SRM IP L and T is blended at ~110 mHz (partially optimized by gradient descent, stopped after 10 iterations, then manually adjusted.) while IP Y is blended at 260 mHz (Lucia's old filter). The stability test was carried over the past weekend with inertial damping and all other stages on. The system seems stable.
  • To study to performance of the inertial damping, I improvised an OL windshield for SRM.
  • Another round of diagonalization was carried just because we have better sensitivity.
  • With IP LVDT damping and other controls on, the residual motion is as follows 9086:

RMS Displacement:

Requirement

Result

L

0.4 µm

0.1076 µm

P

1 µrad

0.0569 µrad

Y

1 µrad

0.0598 µrad

RMS Velocity:

Requirement

Result

L

0.5 µm

0.1014 µm/s

P

/

0.2165 µrad/s

Y

/

0.1776 µrad/s

  • I measured the residual motion with inertial damping, inertial damping without LVDT DC, normal damping without LVDT DC. Results are pretty similar.
  • I modified the optimization algorithm to allow fine learning. The cost plateaued after ~300 iterations and the blending frequency was at 78.6 mHz. I tried implementing this blending filter but the system went unstable.

SR3 Inertial Damping
  • The calibration, diagonalization for IP geophones was done.
  • I measured and fit noise floor for LVDTs and geophones.
  • I found that the IP geophones have a much higher noise level than those from SRM. (1 order of magnitude higher, see 9088)

  • This explains why the optimization result of the blending filter converges much quicker than that of SRM, because the geophone noise level limits the blending frequency to be at ~ 130 mHz.
  • However, when the system went unstable when the blending filter was implemented. I should try a higher blending frequency or manually adjusting the blending filter to have a proper shape.

Type B (Mark)

Report for the week 5/27 (written by Fabian)

  • Shoda-san implemented the moderate watchdog: when tripping the amount of actuation will ramp down to zero smoothly. Klog entry 9009.

  • Terrence updated the BS guardian.
  • Terrence designed a VIS status monitor. The font might be too small for some people.
  • Terrence worked on SRM inertial damping.
    • There is coherence between the geophones and LVDTs at 0.1 Hz.
    • He designed a filter with a blending frequency of 0.110 Hz.
    • The system seems stable.
  • For SR2 Fabian calculated the sensor noise fed back into the system by the control loop: entry 8979.

    • The GAS filters' contrbution is too high.
    • I asked Enzo and he didn't rolled off the filters at higher frequecies.
    • Filter modification was assigned to Fabian.
  • Fabian and Yokozawa-san measured the settings of the SR3 LVDT driver. Klog entry 9025.

  • In the weekly comissioning meeting Fabian made the commitment of writing a list of the measurements we want to do with the DRMI for Type B characterization.I will write it by next week and send it around for comnents.

Type B (Mark)

Report for the week 5/20 (written by Fabian)

  • Terrence implemented a filter for the inertial damping. See entry 8905.

  • The highlight of the week is Terrence's measurement of the residual motion of the SR2 optic: as measured by the oplev, it does fulfill the RMS displacement and velocity requirements for lock acquisition and, partially, for the observation mode also. See the report at klog entry 8949.

    • The integrated RMS was calculated from 5 Hz to DC because there are some peaks between 8 Hz and 10 Hz and whose origin still has to be clarified. See the report at entries 8918 and 8933.

    • After Akutsu-san's suggestion Terrence improved the air shield in two rounds, including a cylindrical cover around the viewport. This reduced the noise by about a factor of 10.
    • He realized setting up the shield may have slightly changed something in the oplev and carried out another round of diagonalization.
    • He realized the H3 coil was not working. We fixed it.A cable was disconnected.
    • The optic residual motion in L showed a large peaks close to 0.2 Hz, which is where the micro-seismic may become apparent. The peak was not seen in P or Y. The inertial damping system is likely not working in a suitable way. The blending frequency is 0.190 Hz which may not be low enough.
    • At the end measurement of the residual motion reported was carried out with inertial damping only in yaw.
  • Fabian, Terrence and Fujii-kun measured the settings of the SR2, BS LVDT drivers and one of SR2's. See klogs 8941 , 8951 and 8952.

  • Fabian worked more on the calculation of the amount of sensor noise injected:
    • I imported Panwei's OSEM noise measurements and calculated the virtual sensor noise.
    • I imported Fujii-kun's measuement of the geophone noise and calculated the virtual sensor noise.
    • I imported the blending filter
    • I wrote lots of comments in Sekiguchi-kun's Matlab scripts.
    • I will write a klog soon with preliminar results.
  • Fabian submitted and abstract for the Autumn JPS meeting on September.

Plan for the week 5/25

  • Terrence will work in the optimization of the inertial damping.
  • Fabian will continue with the calculation and prepare a presentation for the conference in Brazil.

Report for Golden Week

  • Mark monitored vent and pumpdown, and worked on vistools.py.

Report for the week of 5/6

  • Lucia and Terrence got the blending working for the IP LVDTs and geophones.
  • Terrence and Mark modified the models and MEDM screens to add a new IP block, IDAMP, for damping using the blended signal with inertial information, and a new DAMPMODE block for ramping smoothly between IDAMP and the old DAMP.
  • Mark worked on vistools.py, improving the support for cdsMuxMatrix blocks (e.g., OSEM2EUL) and adding support for cdsRampMuxMatrix blocks (e.g., the newly introduced DAMPMODE). He also created two sets of unit tests, unittest.py and unittest.sh, for the two modes of use (as a Python module in other Python programs, and as a command-line utility in bash).

Report for week of 5/13

  • Terrence measured Q's for nearly all modes on SR2 from TFs and then did some ring-down tests to get a better estimate of some high-Q modes.
  • Terrence and Mark looked at the velocity and angular velocity noise of the SR2 as reported by the LEN and TILT OLs.
    • The Y signal was fairly clean and the Y RMS angular velocity requirement was met. However the L signal was quite noisy, and was also contaminating the P signal due to the large cross terms in the diagonalization matrix. The RMS L from 5 Hz down with damping on was around 10 µm/s, which is 10 times the requirement. However little of this was due to pendulum resonances - most of the noise was featureless and broadband. In displacement units the noise was flat up to about 1 Hz and then smoothly rolled off to f^-2. The noise in the raw P and Y signals (not yet divided by the sum) was about 10 times the noise in the SUM.
    • We tried improving the noise with different whitening settings. Increasing the overall gain, increasing it at the QPD but reducing it at the whitening filter, and switching off the 1-10Hz boost whitening filter (and compensating software filter) had no effect. Blocking the beam reduced the noise by about 2 orders and brought the raw P and Y in line with the SUM, suggesting that the problem was before the QPD. Akutsu-san says the noise spectrum looks like it may be limited by air currents, and we probably need better wind shielding.
  • Mark swapped in an improved version of vistools.py, debugged a few small issues, and simplified the TYPEB.py Guardian.

  • Fabian embarked on a project to calculate the residual motion within KAGRA's observation band (beginning at 10 Hz) to make sure The Type B suspension fulfills the requirement, and calculate the speed of the optic to make sure we can acquire the lock.
    • He modified Takanori's Simulink/Matlab simulation in order to include the length sensing oplev and its control feedback path.
    • He imported Enzo's IM/GAS/IP-LVDT filters into the simulation.
    • He imported into the simulation the measured LVDT floor noise (sensitivities).
    • He calculated the IP-L, IP-T and IP-Y sensitivities.

Plan for week of 5/20

  • Mark at Mitaka Mon, Tue; US from Wed till 6/11.
  • Fabian
    • Import Panwei's OSEM sensitivity measurements and calculate the sensitivity of the virtual IM sensors.
    • Import geophones sensitivities and calculate the sensitivity of the corresponding virtual sensors.
    • Import Licia's and Terrence's inertial sensing filters.
  • Terrence
    • Try to reduce the OL noise and get better estimates of the RMS velocity with damping on.

Vacation/Travel Info

  • Terrence's next visit from 4/18 to 7/27.
  • Mark's final day 6/28.

KAGRA/Subgroups/VIS/TypeB/Minutes (last edited 2019-08-16 13:43:37 by fabian.arellano)